A tiny 0.8g ball carries a charge of magnitude 10μC. It is suspended by a thread in a






downward 500N/C electric field. What is the tension in the thread if the charge on the ball is






positive?

1
Expert's answer
2022-05-16T13:40:28-0400
"T=mg+qE\\
= 9.8*0.8*10^{-3}+10*10^{-6}*500\\
=0.0128\:\rm N"